(2b). The average time taken for the nearest Covid-19 ambulance team to convey a patient showing
symptoms to the nearest designated hospital follows the normal distribution about a mean of 60
seconds and a standard deviation of 8 seconds. If 200 patients were selected at random, how many of
them could be taken to the hospital in
i) Less than 50 seconds.
ii) More than 64 seconds.
1
Expert's answer
2020-06-03T19:25:26-0400

Mean =60 seconds

Sd=8 seconds

N=200 patients

i) the patients taken to hospital in less than 50 seconds

"Z=\frac {x-\mu} {\sigma}"

="\frac {50-60}{8}"

=-1.25

P(z<-1.25) is 0.10565

The number of patients out of the 200 that will be taken to the hospital are

=0.10565*200

"\approx 21 patients"

ii) More than 64 seconds

"Z=\frac{64-60}{8}"

=0.5

P(z>0.5)=0.30854 from the standard normal tables

Out of the 200 randomly selected patients

=0.30854*200

"\approx 62 patients" will be taken to the hospital